Proc. 24th Winter Workshop on

Nuclear Dynamics (2008) 000–000  24th Winter Workshop on Nuclear Dynamics South Padre, Texas, USA April 5–12, 2008

Title of the Article (Sometimes Extending over Two or More Lines)

Author One^{1} and Author Two^{2,a}

1 Institute of Author One,

123456 City, Country 2 Institute of Author Two, (if not the same)

456789 City, Country

Abstract. This is a template article, to be used along with the style spadre2008.sty, prepared for manuscripts to the Proceedings of the 24th Winter Workshop on Nuclear Dynamics, South Padre, Texas, April 5–12, 2008. Articles so prepared are suitable for submissions to e.g. the LANL preprint server as well.

Keywords: list of keywords, relevant to the article

PACS: specifications see, e.g. http://www.aip.org/pacs/

1. Introduction

A historical, methodological or logical scope outlined in the Introduction helps the readers place the subject of the article or it enables to guess the present stage of the subject. Ample citation [1–9] of literature that deals with the subject in question is also given in this section, as a source of information. In a second part, the Introduction tells the readers more details about the structure of the article. It often serves as a guide to introduce the sections that build up the article.

At the beginning of the L^{A}TEX template file spadre2008-template.tex, there is a header in which certain data are asked to submit for communication purposes.

2. Technical–Preparatory Section

The technical or methodological preparatory section of the article should commence the row of the sections. The experimental description of measurements or the mathematical introduction of the problem, the definitions of the concepts or the descriptions of the methods, procedures applied in the investigations should be treated in this section to start with the problem [1, 6–9].

Citation of references mentioned in the article will be done as seen in the previous sentence. Citations referring to some journals should be given, for example, as in Ref. [4], using the ‘protocol’ shown in the References.

2.1. Subsections dealing with details

Subsections (maybe, the deeper \subsubsections) will be arranged so that they appear as discernible distinct parts within the whole structure of the article. For too wide tables or figures it may happen that a “landscape page” will be setup.

2.2. Mathematics

Mathematics can be e.g. embedded P_{1}(p) = E_{p}(dN/d^{3}p) as in this case, or displayed separately from text rows as (the boldface letter p denotes a vector):

n_{μ}(x) = ∫_{Σ} d^{3}σ_{μ}(x′) δ^{(4)}(x − x′) .  (1)

Not strange if formulae are arranged in an (vertical or horizontal) array of terms/factors or equations or in matrices, e.g. if a single row is too short to accommodate a formula:

C̃(q, K) = 1 ± {1 − q_{s}^{2}⟨y^{2}⟩ − ⟨[q_{⊥}(x − β_{⊥}t) + q_{L}(z − β_{L}t)]^{2}⟩
+ ⟨q_{⊥}(x − β_{⊥}t) + q_{L}(z − β_{L}t)⟩^{2} + O [⟨(q·x)^{4}⟩]} . (2)

Cross references to formulae will be given as “as in Eq. (2)” or “given by (1)”, etc.

2.3. Tables

An example of a simply built table we give below.

Table 1. Title of the table (usually, the first sentence). An explanatory on the data text may follow; it ranges in the full width of the table, similarly as for figures.

Projectile Energy Ē  σ_{KLL}^{rel}  σ_{KLM}/σ_{KLL}
(MeV) (eV)  (%)  (%)
Ne^{3+}  110  756.6 ± 0.3 71.6 ± 3.8  8.1 ± 1.2
Ne^{10+}  110  742.8 ± 0.2 100  10.3 ± 4.5
Ar^{6+}  220  743.7 ± 0.3 153.5 ± 10.7 < 9.2
Ar^{17+}  220  719.8 ± 0.4 269.2 ± 21.3  9.84 ± 0.35

2.4. Figures

Figure 1 shows a typical presentation of data or results. The figure caption is an inevitable part of the figure (see Fig. 1).

Fig. 1. The figure caption contains an explanatory text for the figure, the meaning of the applied signs, references, explicit data for parameters

Plots are expected to be submitted in Encapsulated PostScript. The epsf.sty style file is used in spadre2008.sty to help you place and size your figures into the manuscript. Please submit each figure as a fig-n.eps(f) file separately. If the file is exceptionally large or in other exceptional cases the compression of files is desirable, however, in this case include instructions for decompressing on a unix computer. Figures will be printed generally in black-&-white in the journals, so please use distinguishable colors, marks and line styles in your figures. Photos should be converted to EPS files; for color separation purposes, please keep the TIF (or good quality JPG) file format if they are colorful photos.

3. Other Sections

Further sections explaining the subject of the article follow the starting ones. At the end, Special Sections may follow as below.

4. Conclusions

As usual, the Conclusions contain the results achieved and presented in the article.

Appendix

Details of mathematical derivations, other technical descriptions are deferred here from the main text. If more than a single appendix will be included, they will be enumerated by using capital Latin letters: Appendix A, Appendix B, etc.

Acknowledgments

A section devoted for personal gratitude and/or offered grant identifications.

Note(s)

(footnotes will be typeset in this way)

a. Permanent address: Institute, City, Country; E-mail: author@host.domain.name

References

1. A.B. Coalman, D.E. Fiedler and G. Hughes, Rev. Mod. Phys. 63 (1991) 545.

2. G.J. Banderer and S.M. Mikhalchuk, Phys. Lett. B54 (1983) 387; Sov. J. Nucl. Phys. 19 (1979) 197.

3. D. Heinemann and U. Sørenmann, Bergensburg preprint BPR-94-31, submitted to Phys. Lett. B. [hep-ph/0123456].

4. T. Czukor and J. Zelényi, Acta Phys. Hung. A 24 (2005) 345.

5. I.V. Mileev, K. Plummer and M.M. Gorentchev, Int. J. Mod. Phys. A11 (1995) 77.

6. S. Trottman and G. Karib, Proceedings of the 5th Workshop on High Energy Collision Phenomena, eds. J.-P. Condrieu and P. Château-Simone, Paris, August 2–6, 1993, Springer, Heidelberg, 1994, p. 324.

7. W.T. Fogg, Talk presented at Advanced Research Workshop on Hot Hadronic Matter: Theory and Experiment, Barcelona, June 29 – July 1, 1994; to be published in the Proceedings by Plenum Press.

8. This was not seen in [6].

9. From Eq. (2) one sees that in the original frame the cross term is given by:

Q_{⊥L}^{2} = (< x τ shη′ > − < x >< τ shη′ >)

up to small logarithmic corrections. This clearly vanishes as soon as the source becomes reflection symmetric under η′ = −η′.